HELP
4) solve the differential equation y+y= èx Satisfying the initial condition Y(0)=1 , Y(o)=-1
Answer #1

PAGE NO: soution. We have, y + y = e i - y lo=1 40) = -1 The auxillary equation associated with is equation homogenous m²+1=0

Please feel free to ask any doubts regarding the solution and please rate positively.

Kind regards.